Can someone please help...
Find an equation that models the path of a satellite if its path is a hyperbola, a = 55,000 km, and c = 81,000 km. Assume that the center of the hyperbola is the origin and the transverse axis is horizontal. Can someone please help me with this.

Answer by stanbon(75887)

Can someone please help...
Find an equation that models the path of a satellite if its path is a hyperbola, a = 55,000 km, and c = 81,000 km. Assume that the center of the hyperbola is the origin and the transverse axis is horizontal.
---------------------
Vertex(55,000,0)
Focus (81,000,0)
----
c^2 = a^2 + b^2
81000^2 = 55,000^2 + b^2
---
b = 59464.27
-------
Form: [x^2/a^2] - [y^2/b^2] = 1
------
x^2/55,000^2 - y^2/59464^2 = 1
=================================
Cheers,
Stan H.